An ornament weighing 36g in air, weighs only 34g in water. Assuming that some copper is mixed with gold to prepare the ornament, find the amount of copper in it. Specific gravity of gold is 19.3and that of copper is 8.9.

Grade:

2 Answers

naveen kumar
11 years ago

mass of ornament in air=36gm  - 1      mass of ornament in water=34gm

desity of gold=19.3gm/cc     density of copper is 8.9gm/cc

loss of wt=wt of displaced water

36-34=2gm

so,  v*p*g=2*g

v=v of gold+v of copper

volume=mass/density

mass of gold/19.3 + mass of copper/8.9 * 1= 2

8.9mass gold + 19.3 mass of copper= 2*19.3*8.9=343.54        - 2

from equations i and 2

m of copper=2.225gm

ankit singh
3 years ago
 

mass of ornament in air=36gm  - 1      mass of ornament in water=34gm

desity of gold=19.3gm/cc     density of copper is 8.9gm/cc

loss of wt=wt of displaced water

36-34=2gm

so,  v*p*g=2*g

v=v of gold+v of copper

volume=mass/density

mass of gold/19.3 + mass of copper/8.9 * 1= 2

8.9mass gold + 19.3 mass of copper= 2*19.3*8.9=343.54        - 2

from equations i and 2

m of copper=2.225gm